Solve each of the following equations for x, and check your results.
26) 7(3x + 4) = 8(2x + 5) + 13

Answer by stanbon(75887)

7(3x + 4) = 8(2x + 5) + 13
21x + 28 = 16x + 40 + 13
21x + 28 = 16x = 53
5x = 25
x=5
----------
Cheers,
Stan H.